OpenStudy (anonymous):

Which of the following is equivalent to log9 36 ?

OpenStudy (anonymous):

9=3^2 36=4*9=2^2*3^2 \[\log_{9} 36=\log_{3^2} (4*9)=\log_{3^2}4 +\log_{3^2} 9=\log_{3^2} 2^2+\log_{3^2} 3^2=\log_{3} 2+1\]

OpenStudy (anonymous):

because \[\log_{x^a} y^b=(b/a)\log_{x} y\]

OpenStudy (anonymous):